OpenStudy (anonymous):

I do not understand this Geometry question. Can some one help? question:Given the triangle and equation below, what is the value of x? Right triangle ABC with AC measuring 7 and BC measuring 6. sec∠C - cot∠A = 3x

OpenStudy (anonymous):

Here is the picture and possible answers:

jimthompson5910 (jim_thompson5910):

Do you know how to find the secant of angle C?

OpenStudy (anonymous):

I think so: Secant of c = 7/6 This is what I did for this part: 1/cosine(x) = 7/6 6 = 7(cosine(x)) .857 = cosine(x) 31.01 = x

jimthompson5910 (jim_thompson5910):

well there's no need to find the angle since we really don't care about the angle itself

jimthompson5910 (jim_thompson5910):

also, it's better leave it as a fraction

jimthompson5910 (jim_thompson5910):

so sec(C) = 7/6

OpenStudy (anonymous):

oh ok

jimthompson5910 (jim_thompson5910):

what about the cotangent of angle A?

OpenStudy (anonymous):

well cotangent = adjacent/opposite but I only have the hypotenuse and the opposite

jimthompson5910 (jim_thompson5910):

do you know how to find the adjacent side?

jimthompson5910 (jim_thompson5910):

Remember that with any right triangle with legs A and B (and hypotenuse C), we know that A^2 + B^2 = C^2

OpenStudy (anonymous):

ya I just thought of that when you posted the question:)

jimthompson5910 (jim_thompson5910):

alright great :)

OpenStudy (anonymous):

so the adjacent side = 3.6

jimthompson5910 (jim_thompson5910):

keep everything as exact as you can

jimthompson5910 (jim_thompson5910):

so keep the square roots

OpenStudy (anonymous):

\[\sqrt{13}\]

jimthompson5910 (jim_thompson5910):

so what did you get before you got 3.6?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so what is cot(A) then?

OpenStudy (anonymous):

so cotangent = \[\sqrt{13}\]/6

OpenStudy (anonymous):

over 6

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so sec∠C - cot∠A = 3x becomes 7/6 - sqrt(13)/6 = 3x

OpenStudy (anonymous):

so just simplify from there

jimthompson5910 (jim_thompson5910):

yes and divide both sides by 3 afterwards to solve for x

OpenStudy (anonymous):

ok so x = 7 - \[\sqrt{13}\] over 18

OpenStudy (anonymous):

That is one of my choices thanks a lot your a great teacher!

jimthompson5910 (jim_thompson5910):

yes I'm assuming you mean \[\Large x = \frac{7-\sqrt{13}}{18}\]

OpenStudy (anonymous):

yup

jimthompson5910 (jim_thompson5910):

thanks, I'm glad I could help you out